# Modeling Quantum Noise in Nanolasers using Markov Chains

### v1

@@ Abstract
The random nature of spontaneous emission leads to unavoidable fluctuations in a laser’s output. This is often included through random Langevin forces in laser rate equations, but this approach falls short for nanolasers. In this paper, we show that the laser quantum noise can be quantitatively computed for a very broad class of lasers by starting from simple and intuitive rate equations and merely assuming that the number of photons and excited electrons only takes discrete values. The success of the model is explained by showing that it constitutes a Markov chain, which can be derived from the full master equations. We show that in the many-photon limit, the model simplifies to Langevin equations. We perform an extensive comparison of different approaches for computing quantum noise in lasers, identifying the best approach for different system sizes, ranging from nanolasers to macroscopic lasers, and different levels of excitation, i.e., cavity photon number. In particular, we find that the numerical solution to the Langevin equations is inaccurate below the laser threshold, while the laser Markov chain model, on the other hand, is accurate for all pump values and laser sizes when collective emitter effects are excluded.
@@ I Introduction
Recently, a stochastic approach to simulating the quantum noise in nanolasers has shown promising results Roy-Choudhury et al. (2009); Roy-Choudhury and Levi (2010); Lebreton et al. (2013); Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025). The approach accurately reproduced photon statistics and laser linewidths calculated using full quantum mechanical master equations (MEs) for microscopic nanolasers with a few emitters Bundgaard-Nielsen et al. (2023, 2025). It also agreed with the well-known Langevin approach in the macroscopic limit of lasers containing thousands of emitters Bundgaard-Nielsen et al. (2025). Importantly, the stochastic approach bridges the mesoscopic regime with ten to hundreds of emitters, where ME calculations become numerically infeasible and Langevin approaches break down due to large noise fluctuations. While alternative methods exist to extend the computational reach beyond direct ME solutions, they all face significant challenges in this regime. Quantum Monte Carlo trajectory methods Mølmer et al. (1993) extend the range of feasibility by evolving a state vector instead of a density matrix, but still suffer exponential scaling. Similarly, the so-called Permutational Invariant Quantum Solver Shammah et al. (2018) reduces the exponential scaling, but as we will see, it does not allow simulations to be extended beyond 10 emitters. Tensor network methods and matrix product states Strathearn et al. (2018); Jørgensen and Pollock (2019); Finsterhölzl et al. (2020); Cygorek and Gauger (2024); Link et al. (2024) reduce the Hilbert space by truncating the eigenvalues of a decomposed state vector, yet remain highly complex and have not yet been extended to nanolasers. Cumulant expansion methods Kubo (1962); Drechsler et al. (2022); Plankensteiner et al. (2022) truncate higher-order quantum correlations in larger systems and are very promising, but require a careful choice of truncation order, and have not been used to compute many-emitter laser linewidths.
@@ II System and Model
Before deriving the stochastic Markov chain model from the ME, we derive the laser rate equations to highlight the essential differences. The typical approach is to derive the equations of motion for the operator expectation values $n_{p}=\expectationvalue{a^{\dagger}a}$ and $n_{e}=\sum_{i}\expectationvalue{\sigma_{i}^{\dagger}\sigma_{i}}$, describing the photon and emitter populations, respectively. This way we obtain ( see e.g., Lorke et al. (2013); Moelbjerg et al. (2013)):
@@
where we defined the total dephasing rate $\gamma_{\perp}=\gamma_{P}+\gamma_{A}+\gamma_{D}+\gamma_{c}$. The problem now is that this equation involves higher-order correlations, such as $\left\langle\sigma_{i}^{\dagger}\sigma_{i}aa^{\dagger}\right\rangle$. To close the otherwise infinite set of equations of motion, we make two approximations. First, the cumulant expansion, where we approximate Fricke (1996); Gartner (2011); Gies et al. (2007); Lorke et al. (2013); Moelbjerg et al. (2013) $\left\langle\sigma_{i}^{\dagger}\sigma_{i}aa^{\dagger}\right\rangle=\left\langle\sigma_{i}^{\dagger}\sigma_{i}\right\rangle\left\langle aa^{\dagger}\right\rangle$ and $\left\langle\sigma_{i}\sigma_{i}^{\dagger}a^{\dagger}a\right\rangle=$ $\left\langle\sigma_{i}\sigma_{i}^{\dagger}\right\rangle\left\langle a^{\dagger}a\right\rangle$. This approximation amounts to ignoring the correlations between the emitter and photon populations; well-justified if $4g<\gamma_{\perp}$. When this approximation holds, we can also take the adiabatic elimination and assume that $d/dt\left\langle\sigma_{i}a^{\dagger}\right\rangle=0$: the emitter polarization follows the photon and emitter populations. This approximation implies that Rabi oscillations and collective effects such as super- and sub-radiant emission processes Auffèves et al. (2011); Gies et al. (2015); Jahnke et al. (2016) cannot be described by the model. With these approximations, we can insert the expression for $\left\langle\sigma_{i}a^{\dagger}\right\rangle$ in Eq. (4) and rework the rate equations to obtain:
@@
We now move on to derive the stochastic LMC from the ME and show that this probabilistic view similarly simplifies to the above rate equation in the many-photon limit. This also means that the LMC can be inferred from the rate equations. This is how the Markov chains have been employed previously Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025), and in Sec. IV.5 we give a concrete example on how to infer the stochastic processes in the Markov chain from eqs. (5)-(6). This gives a simple recipe for translating rate equations to stochastic Markov chains, which do not have the problems of the Langevin approach highlighted in Sec . V.
@@ III Gillespie’s method
The sampling approach can thus be summarized as:
  • Evaluate all $a_{j}(\mathbf{x})$ according to the current state $\mathbf{x}$
  • Generate random values for $\tau$ and $j$
  • Update the time to $t\rightarrow t+\tau$ and population $\mathbf{x}\rightarrow\mathbf{x}+\mathbf{v}_{j}$
  • Record $\mathbf{x}$ and $\tau$ and repeat from step 1 or end the simulation.
This generates a random trajectory of $\mathbf{x}_{j}(t)$ that can be used for subsequent averaging and statistical analysis. To generate samples for $\tau$ and $j$, we choose the First Reaction Method. Here, we draw a random $\tau_{j}$ for each process $j$, which can be done according to:
@@
where $r_{j}$ is a random number equally distributed between 0 and 1. Subsequently, we choose the process $j$ that is associated with the smallest $\tau_{j}$ drawn.
Gillespie’s method is thus a Markov chain that explores the state space $P(\mathbf{x},t)$, which follows an equation of the type in Eq. (7). In the following section, we show that under certain approximations, the ME takes the form of Eq. (7) and thus can be sampled using Gillespie’s method. In Sec. IV.5, we also highlight how one can identify the Markov chain in a simple way directly from a rate equation.
@@ IV.1 Derivation
The off-diagonal elements here play the same role as the polarization $\expectationvalue{\sigma a^{\dagger}}$ plays in the derivation of the laser rate equations in Eq. (4). There (section II), the off-diagonal elements are eliminated using a cumulant expansion and the adiabatic elimination.
In this derivation, the latter step amounts to setting $\partial\rho_{n_{p},g;n_{p}-1,e}/\partial t=0$, to obtain the analytical expression for the equilibrium form of the off-diagonal element $\rho_{n_{p},g;n_{p}-1,e}$; however, we notice the coupling to $\rho_{n_{p}+1,g;n_{p},e}$, in a way similar to how $\expectationvalue{\sigma a^{\dagger}}$ couples to higher order terms. Unlike the cumulant expansion used in section II, here there is no clear interpretation of the role or form of the cumulant expansion. However, assuming:
@@
By extending the previous analysis to each of the $n_{0}$ emitters, we arrive at equations similar to Eqs. (15) and (16) written for each individual emitter. With the introduction of $\rho_{n_{p},n_{e}}$ describing the sum of all density matrix elements with $n_{p}$ number of photons and $n_{e}$ excited emitters and the use of some combinatorics (see appendix B for details), we arrive at:
@@ IV.2 Gillespie’s method applied to the Laser Markov Chain
In the previous section, we derived a Scully-Lamb-like ME, which governs the evolution of the probability distribution of $\rho_{n_{p},n_{e}}$. Defining $\textbf{x}=(n_{p},n_{e})$ and inserting the events from Table 2 into Eq. (7), we recover Eq. (17) line by line. In Appendix B, we explicitly carry out the procedure. Following Gillespie’s method outlined in Section III with the events and rates of Table 2, we can thus efficiently sample the probability distribution outlined in Eq. (17). For a numerical implementation of the scheme, see the supplementary material of Refs. Bundgaard-Nielsen et al. (2023, 2025).
This derivation thus shows a formal connection between the ME and the LMC. This explains the success of the stochastic approach employed in refs. Bundgaard-Nielsen et al. (2023, 2025), which was numerically observed to compare favorably to the ME in contrast with a typical Langevin approach.
In previous works Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025), this formal connection has not been established; instead the events and rates have been inferred from an already derived rate equation. In the following sections, we show how one obtains the rate equations from the LMC in the many-photon limit, thus connecting all the methods for including noise in nanolasers discussed in this paper. We also show that inferring the events and rates from a rate equation is equivalent to the above derivation, which makes the adoption of the LMC methods very simple.
@@ IV.3 Intermediate Photon Limit - Tau-Leaping
If the mean number of photons is large, it can be computationally inefficient to follow all the individual population changes, since several of the rates (see Table 2) scale with the photon number, leading to extremely small time steps. The number of required samples to converge to accurate photon statistics, such as the $g^{(2)}(0)$-function, can thus get very large, since one needs to simulate over time durations that are long compared to the characteristic time scales of the laser, such as the carrier lifetime. Instead, if one assumes that during a time window $\tau$, the rates $a_{j}(\mathbf{x})$ are essentially constant, then the number of times each process $j$ happens is a Poisson random variable with mean $a_{j}(\mathbf{x})\tau$ Gillespie (2007).
@@ IV.4 Many Photon Limit - Langevin Equations
If we can find a $\tau$ where the fractional change $\epsilon$ is small and $a_{j}(\mathbf{x})\tau>>1$, meaning that at each leap many events occur, but they do not change the population significantly, we can approximate the Poisson random variables as Gaussian variables. Using the fact that a gaussian variable with mean $m$ and standard deviation $\sigma$ can be written as $\mathcal{N}(m,\sigma^{2})=m+\sigma\mathcal{N}(0,1)$, we obtain the time evolution:
@@
$D_{aa}$ $=\frac{1}{2}\bigl[\gamma_{r}n_{0}n_{p}+\gamma_{r}n_{e}+\gamma_{c}n_{p}\bigr],$ (23)
$D_{ae}$ $=-\frac{1}{2}\,\gamma_{r}\bigl(n_{e}+n_{0}n_{p}\bigr),$ (24)
@@ IV.5 Laser Markov Chain from Rate Equations
For example, if we consider the rate $\gamma_{r}n_{e}$ which occurs as the second element in Eq. (5) and third element in Eq. (6) with opposite sign, we can immediately identify $\textbf{v}=(+1,-1)$ with the rate $\gamma_{r}n_{e}$. This process is related to spontaneous emission, as shown in Table 2.
The rate $\gamma_{r}(2n_{e}-n_{0})n_{p}$ occurring as first and second term in Eqs.(5)-(6), respectively, is, however, more tricky to interpret. We notice that if $n_{e}<n_{0}/2$ the term changes sign, and we therefore cannot determine the change vector v unambiguously. This hints that there are multiple processes underneath. However, the most obvious separation into $2\gamma_{r}n_{e}$ with the $\textbf{v}=(+1,-1)$ and $n_{0}n_{p}$ with $\textbf{v}=(-1,+1)$ permit unphysical transitions. Specifically, if $n_{e}=n_{0}$, this choice would allow absorption $n_{0}n_{p}$ to take place, resulting in $n_{e}=n_{0}+1$ after the transition - clearly unphysical as we cannot have more excitations than there are emitters present.

### v2

@@ Abstract
The random nature of spontaneous emission leads to unavoidable fluctuations in a laser’s output. This is often included through random Langevin forces in laser rate equations, but this approach falls short for nanolasers. In this paper, we show that the laser quantum noise can be quantitatively computed for a very broad class of lasers by starting from simple and intuitive rate equations and merely assuming that the number of photons and excited electrons only takes discrete values. While the approach has seen previous success, we here derive it rigorously from an open quantum system master equation, whereas it was previously introduced only on phenomenological grounds. We further show that in the many-photon limit, the model simplifies to Langevin equations. We perform an extensive comparison of different approaches for computing quantum noise in lasers, identifying the best approach for different system sizes, ranging from nanolasers to macroscopic lasers, and different levels of excitation, i.e., cavity photon number. In particular, we show that below the laser threshold, stochastic fluctuations in the numerical solution to the Langevin equations can drive populations to unphysical negative values, requiring the introduction of population bounds, which in turn skew the noise statistics, leading to inaccuracies. The Laser Markov Chain model, on the other hand, is accurate for all pump values and laser sizes when collective emitter effects are excluded.
@@ I Introduction
Recently, a stochastic approach to simulating the quantum noise in nanolasers has shown promising results Roy-Choudhury et al. (2009); Roy-Choudhury and Levi (2010); Lebreton et al. (2013); Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025). The approach accurately reproduced photon statistics and laser linewidths calculated using full quantum mechanical master equations (MEs) for microscopic nanolasers with a few emitters Bundgaard-Nielsen et al. (2023, 2025). It also agreed with the well-known Langevin approach in the macroscopic limit of lasers containing thousands of emitters Bundgaard-Nielsen et al. (2025). Importantly, the stochastic approach bridges the mesoscopic regime with ten to hundreds of emitters, where ME calculations become numerically infeasible and Langevin approaches break down due to large noise fluctuations. While alternative methods exist to extend the computational reach beyond direct ME solutions, they all face significant challenges in this regime. Quantum Monte Carlo trajectory methods Mølmer et al. (1993) extend the range of feasibility by evolving a state vector instead of a density matrix, but still suffer exponential scaling. Similarly, the so-called Permutational Invariant Quantum Solver Shammah et al. (2018) reduces the exponential scaling from $\mathcal{O}(4^{N})$ to $\mathcal{O}(N^{4})$ where $N$ is the number of emitters, but as we will see, it does not allow simulations to be extended beyond 10 emitters. Tensor network methods and matrix product states Strathearn et al. (2018); Jørgensen and Pollock (2019); Finsterhölzl et al. (2020); Cygorek and Gauger (2024); Link et al. (2024) reduce the Hilbert space by truncating the eigenvalues of a decomposed state vector, yet remain highly complex and have not yet been extended to nanolasers. Cumulant expansion methods Kubo (1962); Drechsler et al. (2022); Plankensteiner et al. (2022) truncate higher-order quantum correlations in larger systems and are very promising, but require a careful choice of truncation order, and have not been used to compute many-emitter laser linewidths.
@@ II System and Model
Before deriving the Laser Markov Chain model from the ME, we derive the laser rate equations to highlight the essential differences. The typical approach is to derive the equations of motion for the operator expectation values $n_{p}=\expectationvalue{a^{\dagger}a}$ and $n_{e}=\sum_{i}\expectationvalue{\sigma_{i}^{\dagger}\sigma_{i}}$, describing the photon and emitter populations, respectively. This way we obtain ( see e.g., Lorke et al. (2013); Moelbjerg et al. (2013)):
@@
where we defined the total dephasing rate $\gamma_{\perp}=\gamma_{P}+\gamma_{A}+\gamma_{D}+\gamma_{c}$. The problem now is that this equation involves higher-order correlations, such as $\left\langle\sigma_{i}^{\dagger}\sigma_{i}aa^{\dagger}\right\rangle$. To close the otherwise infinite set of equations of motion, we make two approximations. First, the cumulant expansion, where we approximate Fricke (1996); Gartner (2011); Gies et al. (2007); Lorke et al. (2013); Moelbjerg et al. (2013) $\left\langle\sigma_{i}^{\dagger}\sigma_{i}aa^{\dagger}\right\rangle=\left\langle\sigma_{i}^{\dagger}\sigma_{i}\right\rangle\left\langle aa^{\dagger}\right\rangle$ and $\left\langle\sigma_{i}\sigma_{i}^{\dagger}a^{\dagger}a\right\rangle=$ $\left\langle\sigma_{i}\sigma_{i}^{\dagger}\right\rangle\left\langle a^{\dagger}a\right\rangle$. This approximation amounts to ignoring the correlations between the emitter and photon populations; well-justified if $4g<\gamma_{\perp}$. When this approximation holds, we can also take the adiabatic elimination and assume that $d/dt\left\langle\sigma_{i}a^{\dagger}\right\rangle=0$: the emitter polarization follows the photon and emitter populations. This approximation implies that Rabi oscillations and collective effects such as super- and sub-radiant emission processes Maki et al. (1989); Drummond and Raymer (1991); Auffèves et al. (2011); Gies et al. (2015); Jahnke et al. (2016) cannot be described by the model. With these approximations, we can insert the expression for $\left\langle\sigma_{i}a^{\dagger}\right\rangle$ in Eq. (4) and rework the rate equations to obtain:
@@
We now move on to derive the stochastic LMC from the ME and show that this probabilistic view similarly simplifies to the above rate equation in the many-photon limit. This also means that the LMC can be inferred from the rate equations. This is how the Markov chains have been employed previously Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025), and in Sec. IV.5 we give a concrete example on how to infer the stochastic processes in the Markov chain from eqs. (5)-(6). This gives a simple recipe for translating rate equations to Laser Markov Chains, which do not have the problems of the Langevin approach highlighted in Sec . V.
@@ III Gillespie’s method
Gillespie’s method can be summarized as the following numerical algorithm:
  1. 1.
    Initialize the simulation time $t=0$ and the integer-valued population vector $\mathbf{x}(0)=\mathbf{x}_{0}$.
  2. 2.
    For the current state $\mathbf{x}(t)$, evaluate all reaction propensities $a_{j}(\mathbf{x})$ for $j=1,\dots,J$.
  3. 3.
    For each process $j$, draw a random waiting time
@@
    where $r_{j}\in(0,1)$ is uniformly distributed.
  4. 4.
    Identify the process $j^{\ast}$ associated with the smallest waiting time $\tau=\min\{\tau_{j}\}$.
  5. 5.
    Update the simulation time and populations according to the discrete jump
    $t$ $\rightarrow t+\tau,$ (10)
    $\mathbf{x}(t+\tau)$ $=\mathbf{x}(t)+\mathbf{v}_{j^{\ast}},$ (11)
    where $\mathbf{v}_{j^{\ast}}$ is the population update vector for process $j^{\ast}$.
  6. 6.
    Record the updated state $\mathbf{x}(t)$ and repeat from step 2 until the chosen simulation time is reached.
Gillespie’s method is thus a Markov chain that explores the state space $P(\mathbf{x},t)$, which follows an equation of the type in Eq. (7). In the following section, we show that under certain approximations, the ME in Eq. (1) takes the form of Eq. (7) and thus can be sampled using Gillespie’s method. In Sec. IV.5, we also highlight how one can identify the Markov chain in a simple way directly from a rate equation.
@@ IV.1 Derivation
The off-diagonal elements here play the same role as the polarization $\expectationvalue{\sigma a^{\dagger}}$ plays in the derivation of the laser rate equations in Eq. (4). There (section II), the off-diagonal elements are eliminated using a cumulant expansion and the adiabatic elimination $d/dt\left\langle\sigma_{i}a^{\dagger}\right\rangle=0$.
In this derivation, the adiabatic elimination amounts to setting $\partial\rho_{n_{p},g;n_{p}-1,e}/\partial t=0$, to obtain the analytical expression for the equilibrium form of the off-diagonal element $\rho_{n_{p},g;n_{p}-1,e}$; however, we notice the coupling to $\rho_{n_{p}+1,g;n_{p},e}$, in a way similar to how $\expectationvalue{\sigma a^{\dagger}}$ couples to higher order terms. Unlike the cumulant expansion used in section II, here there is no clear interpretation of the role or form of the cumulant expansion. However, assuming:
@@
By extending the previous analysis to each of the $n_{0}$ emitters, we arrive at equations similar to Eqs. (17) and (18) written for each individual emitter. With the introduction of $\rho_{n_{p},n_{e}}$ describing the sum of all density matrix elements with $n_{p}$ number of photons and $n_{e}$ excited emitters and through the use of combinatorics accounting for all the different microscopic emitter configurations (see appendix B for details), we arrive at:
@@ IV.2 Gillespie’s method applied to the Laser Markov Chain
In the previous section, we derived a Scully-Lamb-like ME in Eq. (19), which governs the evolution of the probability distribution of $\rho_{n_{p},n_{e}}$. Since Eq. (19) follows the exact structure of Eq. (7), we can efficiently simulate it using Gillespie’s method. We can further illustrate this by defining $\textbf{x}=(n_{p},n_{e})$ and inserting the events from Table 2 into Eq. (7), see appendix B. We find that we recover Eq. (19) line by line. Following Gillespie’s method outlined in Section III with the events and rates of Table 2, we can thus efficiently sample the probability distribution outlined in Eq. (19). For a numerical implementation of the scheme, see the supplementary material of Refs. Bundgaard-Nielsen et al. (2023, 2025) or Ref.git (2026).
This derivation thus shows a formal connection between the ME in Eq. (1) and the LMC. This explains the success of the stochastic approach employed in refs. Bundgaard-Nielsen et al. (2023, 2025), which was numerically observed to compare favorably to the ME in contrast with a typical Langevin approach.
In previous works Puccioni and Lippi (2015); Mork and Lippi (2018); André et al. (2020); Bundgaard-Nielsen et al. (2023); Takemura et al. (2019); Lippi (2021); Puccioni et al. (2024); Bundgaard-Nielsen et al. (2025), this formal connection had not been established; instead, the events and rates have been inferred from an already derived rate equation. In the following sections, we show how one obtains the rate equations from the LMC in the many-photon limit, thus connecting all the methods for including noise in nanolasers discussed in this paper. We also show that inferring the events and rates from a rate equation is equivalent to the above derivation, which makes the adoption of the LMC methods very simple.
@@ IV.3 Intermediate Photon Limit - Tau-Leaping
If the mean number of photons is large, it can be computationally inefficient to follow all the individual population changes, since several of the rates (see Table 2) scale with the photon number, leading to extremely small time steps. The number of samples required to achieve accurate photon statistics, such as the $g^{(2)}(0)$-function, can thus get very large, since one needs to simulate over time durations that are long compared to the characteristic time scales of the laser, such as the carrier lifetime. Instead, if one assumes that during a time window $\tau$, the rates $a_{j}(\mathbf{x})$ are essentially constant, then the number of times each process $j$ happens is a Poisson random variable with mean $a_{j}(\mathbf{x})\tau$ Gillespie (2007).
@@ IV.4 Many Photon Limit - Langevin Equations
If we can find a $\tau$ for which the fractional change $\epsilon$ is small and $a_{j}(\mathbf{x})\tau>>1$, meaning that at each leap many events occur, but they do not change the population significantly, we can approximate the Poisson random variables as Gaussian variables. Using the fact that a gaussian variable with mean $m$ and standard deviation $\sigma$ can be written as $\mathcal{N}(m,\sigma^{2})=m+\sigma\mathcal{N}(0,1)$, we obtain the time evolution:
@@
$D_{pp}$ $=\frac{1}{2}\bigl[\gamma_{r}n_{0}n_{p}+\gamma_{r}n_{e}+\gamma_{c}n_{p}\bigr],$ (25)
$D_{pe}$ $=-\frac{1}{2}\,\gamma_{r}\bigl(n_{e}+n_{0}n_{p}\bigr),$ (26)
@@ IV.5 Laser Markov Chain from Rate Equations
For example, if we consider the rate $\gamma_{r}n_{e}$, which occurs as the second element in Eq. (5) and the third element in Eq. (6) with opposite sign, we can immediately identify $\textbf{v}=(+1,-1)$ with the rate $\gamma_{r}n_{e}$. This process is related to spontaneous emission, as shown in Table 2.
The rate $\gamma_{r}(2n_{e}-n_{0})n_{p}$, occurring as first and second term in Eqs. (5)-(6), respectively, is, however, more tricky to interpret. We notice that if $n_{e}<n_{0}/2$ the term changes sign, and we therefore cannot determine the change vector v unambiguously. This hints that there are multiple processes underneath. However, the most obvious separation into $2\gamma_{r}n_{e}$ with the $\textbf{v}=(+1,-1)$ and $n_{0}n_{p}$ with $\textbf{v}=(-1,+1)$ permit unphysical transitions. Specifically, if $n_{e}=n_{0}$, this choice would allow absorption $n_{0}n_{p}$ to take place, resulting in $n_{e}=n_{0}+1$ after the transition - clearly unphysical as we cannot have more excitations than there are emitters present.